OpenStudy (anonymous):

i am 4 grade what is 2*34 please

OpenStudy (anonymous):

68

OpenStudy (lncognlto):

2 x 30 = 60 2 x 4 = 8 So 60 + 8 = 68.

HanAkoSolo (jamierox4ev3r):

don't provide the answer, give an explanation @klovic

HanAkoSolo (jamierox4ev3r):

think of the number 34. THen you have a two, so it is basically 34+34, which equals 68

OpenStudy (uri):

Learn your tables..

OpenStudy (goldphenoix):

 What does 4 * 2 give you?

HanAkoSolo (jamierox4ev3r):

lol @uri

hero (hero):

2*34 = 2(30 + 4) = 60 + 8 = 68

OpenStudy (uri):

Do the way @GoldPhenoix said...